# Combinatorial Auctions without a Numeraire: The Case of Blockchain Trade-Intent Auctions

### v1

@@ 5 Equilibrium of the fair combinatorial auction
Finally, the remaining two cases correspond to “uncompetitive batching”: the benefit of specialization is in some intermediate range, and one of the solvers is very productive (relative to the other). This solver can easily disqualify the opponent’s batched bid as unfair. Anticipating no competition from the opponent’s batched bid, this solver wins by bidding the minimal amount to win both orders.
To summarize, when the individual auctions are in first price, each solver can more directly affect the reference for fairness and, as a consequence, will bid higher in the individual auctions. The auction, therefore, provides stronger fairness guarantees because the two traders earn more than $g\underline{\beta}$ and $g\underline{\delta}$, against $\underline{\beta}$ and $\underline{\delta}$ in the simultaneous first-price auctions (c.f., Lemma 1) and in the fair combinatorial auction with individual auctions in second price (c.f., Lemma 3). At the same time, the value of the assets returned to the traders will be lower, especially if the outcome is specialization or “uncompetitive batching”. The comparison with the previous mechanism highlights a trade off in the design of fair combinatorial auctions: providing stronger fairness guarantees comes at the expense of lowering the value of the assets returned to the traders.
6 Discussion.
We now discuss a few ways in which the basic model can be extended.
@@ 6.1 Sequential fair combinatorial auction
An intuitive solution is to have a sequential combinatorial auction: first, solvers bid on the individual trades; then these bids (and the reference outcome for fairness) are revealed; then solvers submit batched bids. The sequential combinatorial auction, however, generates information leakage: how a solver bids in the first phase reveals some information to the opponent, who may use it in the subsequent bidding stage. Because of this, in equilibrium, solvers submit only “uninformative” bids, which is equivalent to skipping the first bidding stage and only bidding on batches, as the next Proposition shows.
@@
One limitation of the above model is that unfair outcomes occur exclusively because there is an indeterminacy regarding how solvers bid in the batch auction. Hence, in a batch auction, next to “unfair” equilibria, there are always also “fair” equilibria. This is an artifact of the model’s simplicity: in this section, we show that if solvers have multiple ways to execute the two orders, then all equilibria of the batch auction may be “unfair”.
To see this, assume that a solver who wins both orders can execute them in two ways. If solver 1 wins both orders, it can choose to generate either $\{g\beta,g\underline{\delta}\}$ (like before) or $\{k\underline{\beta},\tau\underline{\delta}\}$ with $\tau<1$ and $1<g<k$. If, instead, solver 2 wins both orders, it can choose to generate either $\{g\underline{\beta},g\delta\}$ (like before) or $\{k\underline{\beta},\tau\underline{\delta}\}$. We further assume that $k\underline{\beta}>g\overline{\beta}$, so the second option (i.e., $\{k\underline{\beta},\tau\underline{\delta}\}$) always generates more of asset $B$ but less of asset $D$ relative to the other option.
@@
Asset $B$Asset $D$$\tau\cdot\underline{\delta}$$\underline{\delta}$$g\cdot\underline{\delta}$$\underline{\beta}$$\beta$$g\cdot\beta$$k\cdot\underline{\beta}$$g\cdot\beta+p_{DB}\cdot g\cdot\underline{\delta}$$k\cdot\beta+p_{DB}\cdot\tau\cdot\underline{\delta}$
Asset $B$Asset $D$$\tau\cdot\underline{\delta}$$\underline{\delta}$$\delta$$g\cdot\delta$$\underline{\beta}$$g\cdot\underline{\beta}$$k\cdot\underline{\beta}$$g\cdot\beta+p_{DB}\cdot g\cdot\underline{\delta}$$k\cdot\underline{\beta}+p_{DB}\cdot\tau\cdot\underline{\delta}$
@@
Figure 2: Solvers’ production possibilities when matched with both orders (red and green dots) and when matched with a single order (black dot). The shaded areas are the sets of feasible transfers in case a solver wins both orders, as a function of the choice of production.
@@
It turns out that this model variation delivers results identical to those presented earlier. The only difference is that depending on the parameters, in all equilibria of the batch auction, one trader receives less than in the simultaneous standard auctions, even though it is possible to provide both traders with strictly more assets. The equilibrium of the simultaneous standard auctions is unchanged: trader 1 receives $\underline{\beta}$ and trader 2 receives $\underline{\delta}$ (see Lemma 1). The equilibrium of the batch auction is similar to that derived in the main text, but with one difference: if $p_{DB}$ is sufficiently low, then both solvers will produce $\{k\underline{\beta},\tau\underline{\delta}\}$ when matched with both orders. In this case, competition implies that trader 1 receives $k\beta$ and trader 2 receives $\tau\underline{\delta}$. As a consequence, trader 2 receives less than $\underline{\delta}$ in all equilibria of the game, even though it is possible to produce and return tokens in a way that makes both traders better off relative to the two simultaneous standard auctions. If $p_{DB}$ is sufficiently low, the unique equilibrium of the batch auction is, intuitively, unfair.
@@
Concerning the fair combinatorial auction, when the individual-trade auction is in second price, again, the logic discussed earlier continues to hold: there is an equilibrium in which both traders set the individual-trade bids to zero, and the auction does not provide any additional fairness guarantee relative to the batch auction. Finally, when the individual-trade auction is in first price, then the equilibrium is again identical to the one derived earlier: solvers never produce $\{k\underline{\beta},\tau\underline{\delta}\}$ because, if they were to do so, their batched bit would always be disqualified as unfair. Again, both traders earn at least $g\underline{\beta}$ and $g\underline{\delta}$, and the auction is effective at providing some minimal fairness guarantees.
inlineinlinetodo: inlineThe fact that batch auctions are always better than two simultaneous auctions is a strong result. As am extension, we may want to check whether it holds also in a more general version of the model (in the sense of having two arbitrarily type distributions).
@@
6.3 Generalization to multiple traders: fairness as a filter
@@
A natural question is how to extend the above framework to more than two traders and more than two solvers. One way to do so is to have a standard combinatorial auction in which solvers submit multiple batched bids (in addition to individual-trade bids), and then some of these batched bids are filtered out because unfair.
@@ 7 Conclusion
We started by building a theoretical model of trade intent auctions with possible complementarities between trade intents. We then used this model to study the two most common auction formats: (i) batch auctions, in which different trade intents are auctioned off jointly as a batch; and (ii) simultaneous standard auctions, in which there is a separate auction for each trade intent (of which Dutch auctions are a subcase). The main result of this comparison is that batch auctions generate more competition between solvers than simultaneous standard auctions. The reason is that batching forces all solvers to compete: even in cases where there is a “best” solver at each trade, the batch auction forces each “best” solver to outbid the other to win any trade. At least under the assumptions of our model, we find that batch auctions always deliver higher overall value to the traders than individual trade-intent auctions (such as Dutch auctions). At the same time, in a batch auction, the solvers’ payoff only depends on the value of the assets returned to the traders, not how this value is shared. Therefore, the benefit of the batch auction may accrue disproportionally to one trader, leaving the other worse off relative to the simultaneous standard auctions.
We then use the model to study a new type of auction: the fair combinatorial auction. This type of auction can be reduced to a combination of batch auctions and simultaneous standard auctions, because solvers can submit bids on individual trades and also on batches of trades. The key assumption is that the outcome of the individual trade-intent auctions (as constructed using the individual-trade bids) is used to filter out batched bids: batched bids are considered by the auctioneer only when they are better than the outcome of the individual trade auctions for all traders in the batch. The outcome of the individual trade auctions can, therefore, be seen as a benchmark for fairness: all batched solutions that improve upon this outcome for all traders are “fair” and considered in the auction, and all other batched solutions are discarded as “unfair”.
@@
We illustrates this point by showing that when the individual trade-intent auctions are second-price auctions, the fair combinatorial auction does not provide any additional fairness guarantees relative to the simple batch auction. The reason is that there are equilibria in which all solvers bid the minimum amount on individual trades, which is equivalent to skipping the individual auctions and only having a batch auction. If the individual auctions are in second price, then no solver can change this outcome by changing its bid on the individual auctions. Instead, when the individual trade-intent auctions are first-price auctions, each solver can manipulate the reference for fairness with its bids on individual trades, independently of the bids by other solvers. In this case, we show that the fair combinatorial auction guarantees that all traders receive more than what they would earn in the simultaneous standard auctions. Hence, even though not all fair combinatorial auctions provide strong fairness guarantees in equilibrium, some do. However, these higher fairness guarantees come at the expense of lower value returned to the traders.
@@
To show that $Q^{*}_{2}(B)=g\underline{\beta}$ and $Q^{*}_{1}(D)=g\underline{\delta}$, we proceed by contradiction: suppose that $Q^{*}_{2}(B)<g\underline{\beta}$. The fact that the total value of the batched bid must be greater or equal to $g\underline{\beta}+p_{BD}g\underline{\delta}$ implies that $Q^{*}_{2}(D)>g\delta$. Hence, the probability that solver 2’s batched bid is considered unfair because of order 2 is zero, while this probability is strictly positive for order 1. Solver 2 is better off by decreasing $Q^{*}_{2}(D)$ and increasing $Q^{*}_{2}(B)$ to keep the total value of the batched bid constant because doing so reduces the probability that the batched bid is considered as unfair. Hence, it must be that $Q^{*}_{2}(B)=g\cdot\underline{\beta},\leavevmode\nobreak\ Q^{*}_{2}(D)<g\delta$. The same reasoning can be applied to show that $Q^{*}_{1}(B)<g\beta,\leavevmode\nobreak\ Q^{*}_{1}(D)=g\cdot\underline{\delta}$.
@@
First, we prove that, in any equilibrium, solvers bid fixed amounts in the first bidding stage. We do so by contradiction. Suppose that, in equilibrium, solver 1’s first stage bid is an increasing function of $\beta$ (note: it could be globally increasing or locally increasing). If this is the case, upon observing a first-period bid from solver 1, solver 2 can infer the maximum batched bid that solver 1 can submit and try to outbid solver 1 by a minimal amount. Anticipating this, in the first period of the game, solver 1 should always bid the minimal amount to deceive solver 2 into submitting a low batched bid. Hence, solver 1’s submitting a first-stage bid is an increasing function of $\beta$ is not an equilibrium. An identical argument rules out equilibria in which solver 2’s submits a first-stage bid that is an increasing function of $\delta$. inlineinlinetodo: inlinethe proof shows that there cannot be a pure strategy equilibrium, but I believe the argument extends to mixed strategy equilibria: in these equilibria, the realized bid does not exactly reveal the type, but does allow the opponent to make some inference and hence adjust the batched bid

### v2

@@ 5 Equilibrium of the fair combinatorial auction
Finally, the remaining two cases correspond to “uncompetitive batching”: the benefit of specialization is in some intermediate range, and one of the solvers is very productive (relative to the other). This solver can easily disqualify the opponent’s batched bid as unfair. Anticipating no competition from the opponent’s batched bid, this solver wins by bidding the minimal amount required for its batched bid to be considered fair.
To summarize, when the individual auctions are in first price, each solver can more directly affect the reference for fairness. In equilibrium, each solver bids higher than in the individual simultaneous auctions. The auction, therefore, provides strong fairness guarantees because the two traders earn at least $g\underline{\beta}$ and $g\underline{\delta}$ (against $\underline{\beta}$ and $\underline{\delta}$ in the simultaneous first-price auctions, c.f., Lemma 1). At the same time, the value of the assets returned to the traders will be lower than that of a batch auction, especially if the outcome is specialization or “uncompetitive batching.”. The comparison with the previous mechanism highlights a trade off in the design of fair combinatorial auctions: providing stronger fairness guarantees comes at the expense of lowering the value of the assets returned to the traders.
Finally, note that the trade-off between fairness and value returned to the traders depends on $g$. If $g$ is sufficiently large, the equilibrium is competitive batching, and there is no trade-off: the value returned to the traders in the fair combinatorial auction is equal to that of the batch auction, and the outcome is fair. The lower is $g$, instead, the less value the auction returns to the user (while maintaining the fairness guarantees). Intuitively, when the benefit of batching is large, it is easier to share those benefits in a way that is “fair” to both traders, while when these benefits are low, being fair requires distorting the outcome of the auction (relative to the simple batch auction).
6 Discussions and extensions.
We now discuss some features of the model and a few ways in which it can be extended.
6.1 Benchmark with frictionless financial markets
It may be useful to move away from our main assumption and consider how the problem changes if financial markets are perfect, meaning if anyone can frictionlessly exchange financial assets at their commonly-known market prices.
Call $p_{BA}$ the market price of asset B in terms of asset A, and $p_{DC}$ the market price of asset D in terms of asset C. If financial markets are perfect, then the equilibrium of the simultaneous standard auctions must be $p_{BA}$ and $p_{DC}$: both solvers can return at least $p_{BA}$ and $p_{DC}$ by accessing the financial market, and no solver will want to return more because they would rather sell the excess on the financial markets. But because $p_{BA}$ and $p_{DC}$ are common knowledge, then the auction designer knows the equilibrium of the simultaneous standard auctions. To prevent the emergence of “unfair” equilibria in the batch auction, it is enough to introduce two limit prices equal to $p_{BA}$ and $p_{DC}$.
To summarize, the fact that trading on financial markets is subject to frictions and fees is a necessary condition for the problem of fairness to be relevant. The reason is that the auction designer does not know these frictions and fees and hence cannot determine the equilibrium of the simultaneous standard auctions. The reference for fairness needs to be elicited as part of the mechanism.
@@ 6.2 Sequential fair combinatorial auction
An intuitive solution is to have a sequential combinatorial auction: first, solvers bid on the individual trades; then these bids (and the reference outcome for fairness) are revealed; then solvers submit batched bids. The sequential combinatorial auction, however, generates information leakage: how a solver bids in the first phase reveals some information to its opponent, who may use it in the subsequent bidding stage. Because of this, in equilibrium, solvers submit only “uninformative” bids in the first stage, making the equilibrium of the game identical to that of a standard batch auction, as the next Proposition shows.
@@
One limitation of the above model is that unfair outcomes occur exclusively because there is an indeterminacy regarding how solvers bid in the batch auction. Hence, in a batch auction, next to “unfair” equilibria, there are always also “fair” equilibria. This is an artifact of the model’s simplicity: in this section, we show that if solvers have multiple ways to execute the two orders, then it is possible that all equilibria of the batch auction are “unfair”.
Assume that a solver who wins both orders can execute them in two ways. If solver 1 wins both orders, it can choose to generate either $\{g\beta,g\underline{\delta}\}$ (like before) or $\{k\underline{\beta},\tau\underline{\delta}\}$ with $\tau<1$ and $1<g<k$. If, instead, solver 2 wins both orders, it can choose to generate either $\{g\underline{\beta},g\delta\}$ (like before) or $\{k\underline{\beta},\tau\underline{\delta}\}$. We further assume that $k\underline{\beta}>g\overline{\beta}$, so the second option (i.e., $\{k\underline{\beta},\tau\underline{\delta}\}$) always generates more of asset $B$ but less of asset $D$ relative to the other option.
@@
Asset $B$Asset $D$$\tau\underline{\delta}$$\underline{\delta}$$g\underline{\delta}$$\underline{\beta}$$\beta$$g\beta$$k\underline{\beta}$$g\cdot\beta+p_{DB}\cdot g\cdot\underline{\delta}$$k\cdot\beta+p_{DB}\cdot\tau\cdot\underline{\delta}$sset $B$Asset $D$$\tau\underline{\delta}$$\underline{\delta}$$g\underline{\delta}$$g\cdot\beta+p_{DB}\cdot g\cdot\underline{\delta}$$k\cdot\beta+p_{DB}\cdot\tau\cdot\underline{\delta}$
Asset $B$Asset $D$$\tau\underline{\delta}$$\underline{\delta}$$\delta$$g\cdot\delta$$\underline{\beta}$$g\underline{\beta}$$k\underline{\beta}$$g\cdot\beta+p_{DB}\cdot g\cdot\underline{\delta}$$k\cdot\underline{\beta}+p_{DB}\cdot\tau\cdot\underline{\delta}$sset $B$Asset $D$$\tau\underline{\delta}$$\underline{\delta}$$\delta$$g\delta$$g\cdot\beta+p_{DB}\cdot g\cdot\underline{\delta}$$k\cdot\underline{\beta}+p_{DB}\cdot\tau\cdot\underline{\delta}$
@@
Figure 2: Solvers’ production possibilities when matched with both orders (red and green dots) and when matched with a single order (black dot). The shaded areas are the sets of feasible transfers in case a solver wins both orders, as a function of the choice of production. On the top, $p_{DB}$ is low and $\{k\cdot\underline{\beta},\tau\underline{\delta}\}$ generates higher value than the alternative; on the bottom $p_{DB}$ is high and $\{k\cdot\underline{\beta},\tau\underline{\delta}\}$ generates lower value than the alternative.
@@
In this variation of the model, the equilibrium of the simultaneous standard auctions is identical to that derived earlier: trader 1 receives $\underline{\beta}$ and trader 2 receives $\underline{\delta}$ (see Lemma 1). The equilibrium of the batch auction is similar to that derived in the main text, but with one difference: if $p_{DB}$ is sufficiently low, then both solvers will produce $\{k\underline{\beta},\tau\underline{\delta}\}$ when matched with both orders. In this case, competition implies that trader 1 receives $k\underline{\beta}$ and trader 2 receives $\tau\underline{\delta}$. Trader 2 receives less than $\underline{\delta}$ in all equilibria of the game, even though it is possible to produce and return tokens in a way that makes both traders better off relative to the two simultaneous standard auctions. If $p_{DB}$ is sufficiently low, the unique equilibrium of the batch auction is, intuitively, unfair.
@@
Concerning the fair combinatorial auction, when the individual-trade auction is in second price, again, the logic discussed earlier continues to hold: there is an equilibrium in which both traders set the individual-trade bids to zero, and the auction does not provide any additional fairness guarantee relative to the batch auction. If, instead, the individual-trade auction is in first price, then the equilibrium is again identical to the one derived earlier: solvers never produce $\{k\underline{\beta},\tau\underline{\delta}\}$ because, if they were to do so, their batched bit would always be disqualified as unfair. Again, both traders earn at least $g\underline{\beta}$ and $g\underline{\delta}$, and the auction is effective at providing some minimal fairness guarantees.
@@
6.4 Generalization of the fair combinatorial auction to multiple traders: fairness as a filter
@@
A natural question is how to define the fair combinatorial auction whenever there are more than two traders and more than two solvers. One way to do so is to have a combinatorial auction in which, first, solvers submit multiple batched bids (in addition to individual-trade bids), then some of these batched bids are filtered out because they are unfair, and finally, the auctioneer selects the combination of surviving bids that maximizing the value of tokens returned to the traders.
@@
6.5 Are batch auctions always better than simultaneous standard auctions?
In our model, the batch auction generates more competition between solves and hence always returns more value to users than the simultaneous standard auctions. Intuitively, in a batch auction, each solver competes with all other solvers. However, if solvers are specialized, then in each simultaneous auction, each solver competes only with the other solvers specializing in that type of trade. This result can be easily extended. For example, suppose there are multiple solvers, multiple types of trades, and multiple trades of each type. If, again, each solver is specialized in a single type of trade, then the simultaneous standard auctions will be uncompetitive, and the specialized solver can win the auction by bidding what any other solver could have produced. Again, a batch auction generates more competition between solvers and more value for the traders.
This result, however, does not extend to all possible variations of the model. For example, suppose there are two identical solvers for each type of trade. In this case, each simultaneous standard auction will be very competitive, with the two identical solvers competing for the right to fill the order. Although batch auctions generate more competition than simultaneous standard auctions, the level of competition in the simultaneous standard auctions will nonetheless be high. If $g$ is sufficiently low, the simultaneous standard auctions may generate more value than the batch auction.
To summarize, the important observation is that batch auctions generate more competition than simultaneous standard auctions. Whether this results in more value being returned by one type of auction depends on the level of competition of the simultaneous standard auctions and the benefit of batching vs the benefit of specialization. We conclude by noting that in the fair combinatorial auction, each solver competes both with other solvers specialized in the same type of trades (via the individual-trade bids) and also with all solvers’ batched bids.
@@ 7 Conclusion
We started by building a theoretical model of trade intent auctions. We then used this model to study the two most common auction formats: (i) batch auctions, in which different trade intents are auctioned off jointly as a batch; and (ii) simultaneous standard auctions, in which there is a separate auction for each trade intent (of which Dutch auctions are a subcase). The main result of this comparison is that batch auctions generate more competition between solvers than simultaneous standard auctions. The reason is that batching forces all solvers to compete: even in cases where there is a “best” solver at each trade, the batch auction forces each “best” solver to outbid the other to win any trade. At least under the assumptions of our model, we find that batch auctions always deliver higher overall value to the traders than individual trade-intent auctions (such as Dutch auctions). At the same time, in a batch auction, the solvers’ payoff only depends on the value of the assets returned to the traders, not how this value is shared. Therefore, the benefit of the batch auction may accrue disproportionally to one trader, leaving the other worse off relative to the simultaneous standard auctions.
Motivated by this observation, we then introduce a new type of auction: the fair combinatorial auction. This type of auction can be reduced to a combination of batch auctions and simultaneous standard auctions, because solvers can submit bids on individual trades and also on batches of trades. The key assumption is that the outcome of the individual trade-intent auctions (as constructed using the individual-trade bids) is used to filter out batched bids: batched bids are considered by the auctioneer only when they are better than the outcome of the individual trade auctions for all traders in the batch. The outcome of the individual trade auctions can, therefore, be seen as a benchmark for fairness: all batched solutions that improve upon this outcome for all traders are “fair” and considered in the auction, and all other batched solutions are discarded as “unfair”.
@@
We illustrate this point by showing that when the individual trade-intent auctions are second-price auctions, the fair combinatorial auction does not provide any additional fairness guarantees relative to the simple batch auction. In this auction, there are equilibria in which all solvers bid the minimum amount on individual trades, because no solver can change this outcome by changing its bid on the individual auctions. Such equilibria are equivalent to skipping the individual auctions and only having a batch auction. Instead, when the individual trade-intent auctions are first-price auctions, each solver can manipulate the reference for fairness with its bids on individual trades, independently of the bids by other solvers. In this case, we show that the fair combinatorial auction guarantees that all traders receive more than what they would earn in the simultaneous standard auctions. Hence, even though not all fair combinatorial auctions provide strong fairness guarantees in equilibrium, some do. However, these higher fairness guarantees come at the expense of lower value returned to the traders.
@@
The first step of the proof is to show that if $g>1$, then in any equilibrium of the auction, $Q^{*}_{1}(B)>\underline{\beta}$ and $Q^{*}_{2}(D)>\underline{\delta}$. We show this by contradiction: suppose there is an equilibrium with $Q^{*}_{1}(B)\leq\underline{\beta}$ (the case $Q^{*}_{2}(D)\leq\underline{\delta}$ is analogous). Then, solver 1’s batched bid is always considered unfair. In equilibrium, if solver 2’s batched bid is fair, then solver 1 wins no trade; if solver 2’s batched bid is unfair, then solver 1 wins only order 1. However, if solver 1 deviates and sets $Q_{1}(B)>q^{*}_{1}(B)\geq\underline{\beta}$ and $Q_{2}=g\cdot\underline{\delta}$, then with strictly positive probability, it wins both orders. As long as $Q_{1}(B)<g\beta$, then this deviation is profitable, and there is no equilibrium with $Q^{*}_{1}(B)\leq\underline{\beta}$.
@@
We now show that, if $g>1$, then $Q^{*}_{2}(B)=g\underline{\beta}$ and $Q^{*}_{1}(D)=g\underline{\delta}$. We proceed again by contradiction: suppose that $Q^{*}_{2}(B)<g\underline{\beta}$. The fact that $Q^{*}_{2}(D)>\underline{\delta}$ implies that the probability that solver 2’s batched bid is considered unfair because order 2 is zero, while this probability is strictly positive for order 1. Solver 2 is better off by decreasing $Q^{*}_{2}(D)$ and increasing $Q^{*}_{2}(B)$ to keep the total value of the batched bid constant because doing so reduces the probability that the batched bid is considered as unfair. Hence, it must be that $Q^{*}_{2}(B)=g\cdot\underline{\beta},\leavevmode\nobreak\ Q^{*}_{2}(D)<g\delta$. The same reasoning can be applied to show that $Q^{*}_{1}(B)<g\beta,\leavevmode\nobreak\ Q^{*}_{1}(D)=g\cdot\underline{\delta}$.
Finally, in case $g=1$, it is easy to see that any feasible batched bid is always considered unfair. The equilibrium is, therefore, identical to the simultaneous first-price auctions. Any $Q^{*}_{1}(B)\leq\underline{\beta}$, $Q^{*}_{2}(D)\leq\underline{\delta}$ is an equilibrium, including $Q^{*}_{1}(B)=\underline{\beta}=g\underline{\beta}$, $Q^{*}_{2}(D)=\underline{\delta}=g\underline{\beta}$. Hence, whereas when $g>1$, then $Q^{*}_{2}(B)=g\underline{\beta}$ and $Q^{*}_{1}(D)=g\underline{\delta}$ in all equilibria of the game, when $g=1$ focusing on the case $Q^{*}_{2}(B)=g\underline{\beta}$ and $Q^{*}_{1}(D)=g\underline{\delta}$ is without loss of generality.
Proof of Lemma 5.
Write (4) and (5) as
$\begin{cases}X_{1}\equiv(\beta-g\underline{\beta})\,\mathrm{pr}\{q_{2}(D)\geq g\underline{\delta}\}+(\beta\cdot g-g\underline{\beta})\,\mathrm{pr}\{q_{2}(D)<g\underline{\delta}\}&\text{if }q_{1}(B)=Q_{1}(B)=g\underline{\beta}\\ \\ \begin{aligned} Y_{1}\equiv&(\beta\cdot g-\tilde{Q}_{1}(B))\,\mathrm{pr}\{q_{2}(D)<g\underline{\delta}\}\\ &\quad\quad\times\,\mathrm{pr}\{\tilde{Q}_{1}(B)>p_{DB}Q_{2}(D)+V\mid q_{2}(D)<g\underline{\delta}\}\end{aligned}&\text{if }q_{1}(B)=\underline{\beta},Q_{1}(B)=\tilde{Q}_{1}(B)\end{cases}$
$\begin{cases}X_{2}\equiv p_{DB}\left((\delta-g\underline{\delta})\,\mathrm{pr}\{q_{1}(B)\geq g\underline{\beta}\}+(\delta\cdot g-g\underline{\delta})\,\mathrm{pr}\{q_{1}(B)<g\underline{\beta}\}\right)&\text{if }q_{2}(D)=Q_{2}(D)=g\underline{\delta}\\ \\ \begin{aligned} Y_{2}\equiv&p_{DB}\left((\delta\cdot g-\tilde{Q}_{2}(D))\,\mathrm{pr}\{q_{1}(B)<g\underline{\beta}\}\right.\\ &\left.\quad\times\,\mathrm{pr}\{p_{DB}Q_{2}(D)+V>Q_{1}(B)\mid q_{1}(B)<g\underline{\beta}\}\right)\end{aligned}&\text{if }q_{1}(B)=\underline{\beta},Q_{1}(B)=\tilde{Q}_{1}(B)\end{cases}$
We want to show that in every equilibrium of the game $X_{1}>Y_{1}$ and $X_{2}>Y_{2}$. We proceed by contradiction.
Suppose first that $X_{1}\leq Y_{1}$ and $X_{2}\leq Y_{2}$. In this case, the equilibrium is identical to a standard batch auction, but for the fact that $Q^{*}_{1}(D)$ and $Q^{*}_{2}(B)$ are pinned down by Lemma 4. The derivations in Section 3.2 (in particular, Lemma 2 and the subsequent discussion) show that, for every realization of $\beta$ and $\delta$, in equilibrium, each solver submits a bid of value at least as large as the minimum value that each solver can generate, that is
$Q^{*}_{1}(B)+p_{DB}Q^{*}_{1}(D)\geq g\cdot\underline{\beta}+p_{DB}\cdot g\cdot\underline{\delta}$
and
$Q^{*}_{2}(B)+p_{DB}Q^{*}_{2}(D)\geq g\cdot\underline{\beta}+p_{DB}\cdot g\cdot\underline{\delta}.$
But together with Lemma 4, the above conditions imply $Q^{*}_{1}(B)\geq g\cdot\underline{\beta}$ and $Q^{*}_{2}(D)\geq g\cdot\underline{\delta}$, which then imply $X_{1}>Y_{1}$ and $X_{2}>Y_{2}$, therefore contradicting our initial hypothesis.
Suppose now that $X_{1}\leq Y_{1}$ and $X_{2}>Y_{2}$ (the argument is identical for the case $X_{1}>Y_{1}$ and $X_{2}\leq Y_{2}$). If $\delta\geq g\underline{\delta}$, then solver 2 sets $q_{2}(D)=g\underline{\delta}$ and wins both trades. If $\delta<g\underline{\delta}$, then the equilibrium is identical to a batch auction, but for the fact that solver 2’s type distribution is truncated from above. Nonetheless, also in this case, in equilibrium, both solvers return more than the minimum value they generate. Again, this implies that $Q^{*}_{1}(B)\geq g\cdot\underline{\beta}$ and $Q^{*}_{2}(D)\geq g\cdot\underline{\delta}$, which then imply $X_{1}>Y_{1}$ and $X_{2}>Y_{2}$, therefore contradicting our initial hypothesis.
@@
First, we prove that, in any equilibrium, solvers bid fixed amounts in the first bidding stage. We do so by contradiction. Suppose that, in equilibrium, solver 1’s first stage bid is an increasing function of $\beta$ (note: it could be globally increasing or locally increasing). If this is the case, upon observing a first-period bid from solver 1, solver 2 can infer the maximum batched bid that solver 1 can submit and try to outbid solver 1 by a minimal amount. Anticipating this, in the first period of the game, solver 1 should always bid the minimal amount to deceive solver 2 into submitting a low batched bid. Hence, solver 1’s submitting a first-stage bid is an increasing function of $\beta$ is not an equilibrium. An identical argument rules out equilibria in which solver 2’s submits a first-stage bid that is an increasing function of $\delta$.